# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

### v2

@@ 6 Evaluation
One of the challenges in model evaluation is handling irrelevant predictions from models. Even if models follow instructions and produce parts of the expected solution, we regularly observe that they — especially open-source models — append a continuation of task irrelevant predictions. Such predictions result in failure modes when applying conditions and validations, and halt any multi-step procedure. Our solution is an evaluation protocol that refines the performance measurement, allowing for more nuanced diagnostics and the possibility of continuing the evaluation despite intermediate failures. To derive our quality measure, we borrow ideas from the utilization of the Fréchet distance for generative processes (Heusel et al., 2017).
@@
We generate trajectories through a NeSy sequential process that creates a trajectory of distributions $\mathbb{P}$ over multiple iterations of generative nodes. Each node in the process can be aligned to a reference distribution, which marks the desired behavior. To quantify the validity of the generated trajectories, we measure the total distance between the generated and reference data distribution along the path trajectory. We therefore adopt a cumulative measure capable of taking into account the entire generative trajectory. In theory, this process would entail calculating the path integral over the latent space representations for models, cumulating the Fréchet distances (Dowson & Landau, 1982) traversed along these distributional trajectories:
@@
However, this approach is computationally intractable for large-scale problems, and requires access to latent representations, which — especially in the context of LLMs — is not always given. For computational feasibility, we therefore introduce an approximation that measures the embedding distances over the path trajectories utilizing an auxiliary embedding model, based on prior work on distribution regression (Szabó et al., 2016). The embedding model maps the symbolic representations into a RKHS, such that we can apply a kernel mean embedding function to measure their respective distances (You et al., 2019; Dinu et al., 2023). We assess the distance through the mean embeddings w.r.t. to a kernel function $K(\cdot,\cdot)$ of the samples $\mathbf{e}^{t}_{x}\sim\nu^{t}_{\text{gen}}\in\mathbb{P}_{\text{gen}}$ and $\mathbf{e}^{t}_{y}\sim\nu^{t}_{\text{ref}}\in\mathbb{P}_{\text{ref}}$ produced by the generated data distribution and a reference data distribution respectively. We denote by $\mu_{\mathbf{e}^{t}_{x}}$, $\mu_{\mathbf{e}^{t}_{y}}$ the mean embeddings associated to the respective samples, i.e. $\mu_{\mathbf{e}^{t}_{x}}(z)=\frac{1}{n}\sum_{i=1}^{n}K(x_{i}^{t},z)$ in case $\mathbf{e}^{t}_{x}=(x_{i}^{t})_{i=1}^{n}$ is a sample of size $n$ of the respective mean embeddings. To compute the similarity between the embeddings of the generated and reference distributions, we evaluate the associated maximum mean discrepancy $\text{MMD}^{2}(\mu_{\mathbf{e}^{t}_{x}},\mu_{\mathbf{e}^{t}_{y}})$ (Gretton et al., 2012) and then, as before for the Fréchet distances, we integrate over $t$:
@@
In empirical evaluations, however, we care about normalized values for ease of interpretation. We therefore analyze the properties of the MMD and derive a similarity score, which follows the same statistical principles as the MMD, and is bound between $[0,1]$. We concluded that we can use only the MMD cross terms to evaluate the similarities. See our supplementary materials Section A for more details. For our comparisons as referenced in Figure 5 we therefore denote the similarities rather than distances. We then come to the following formulation and refer to our empirical measure as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short:
@@
We approximate the integral across time steps through Monte Carlo approximation. The introduced normalization constants denote the similarities to a random sequence $z_{\text{rand}}$, which functions as a baseline subtraction to recenter our results, and a given reference score to rescale w.r.t. to scores obtained from comparing related solutions $z$. The use of the min-max scaling ensures the final measure is bounded between $[0,1]$. This process reflects properties such as Hölder continuity that bounds the kernel function within certain limits, ensuring that the similarity measure is appropriately scaled. To compute the embeddings, we utilize the embedding model all-mpnet-base-v2 (Song et al., 2020), due to its widespread availability, and its balance between speed and quality. As a similarity measure, we select a Gaussian kernel following our derivation from the supplementary materials Section A. In our supplementary implementations, we also explore other kernels, including some preliminary experiments with cosine similarity. We also note that one can integrate Bernoulli distributed trials into our score, with 00 values representing failure modes and values of $1$ being successes. Furthermore, if we relax our definition, we can integrate other similarity measures which are bound between $[0,1]$, which then reflect on domain-specific attributions, i.e. including a similarity measure tailored towards capturing the nuances between two sub-structures of abstract syntax tree.
In our evaluation of Figure 5 we conclude with the cumulative score, for the following base performance criteria.
Basic Capabilities
We analyze the consistency and stability of models and establish baselines by addressing the following performance criteria:
  • Associative Prediction: We evaluate the success rate of models to follow simple and complex instructions and associations with zero- and few-shot examples. We therefore address the proficient use of our operators between Symbol types.
  • Multi-modal Binding: We perform data transformations between multiple modalities by binding through language-based representations, and evaluate their proficiency in tool utilization, classification and routing of requests to relevant modules.
  • Program Synthesis: We evaluate executable code with and without including concepts from retrieval augmented generation, model-driven development, such as templating to direct the generative flow, and experiment with self-generated instructions by creating self-referential expressions. We not only address similarity between produced solutions, but also include the measurement of abstract syntax tree equivalence and the success of executions.
@@
To evaluation the logic capabilities of models, we conditioned them to create a sequence of expressions as self-contained components, and refer to higher-order predicate logic for their assessment. Based on the underlying type theory originating from Whitehead & Russell (1925–1927), we evaluate a model’s capability in the form of there exists x such that x fulfills y. Such quantifiers define the standard semantics over expressions, where the meaning is given by a semantic function. A semantic function maps a term from an abstract definition to a point in a domain, which is an interpretation of the term’s type and value. Therefore, these functions operate on types and values of expressions, and relations thereof. Subsequently, NeSy engines can formulate and evaluate at inference time logic-based instructions through Lisp, Prolog, or Mathematica (McCarthy, 1959; Colmerauer & Roussel, 1993; Chen et al., 1993; Inc., 2022), or leverage SAT solvers such as Z3 (Moura & Bjørner, 2008). Therefore, the evaluation of a natural language statement can be interpreted by any expert system which defines the corresponding semantic functions and process them either in a symbolic (Feigenbaum et al., 1965; Gamble et al., 1994), differentiable (Veličković & Blundell, 2021; Ibarz et al., 2022), or hybrid manner (Kuncicky et al., 1991).
We evaluate how proficient models interpret custom DSLs and define expression statements. The DSL is designed to express logical relations and operations in a structured format, and supports human-readable and machine-interpretable formulations. The following example illustrates such relationships by translating a natural language statement into an expression statement, as follows:
@@
A DSL may enforce the usage of $\text{HAS}(\cdot)$, $\text{IS}(\cdot)$, etc. and may condition a LLM to produce the following expressions:
@@
These are then used to define the following logical expression:
@@
An automated theorem prover can now evaluate this statement for all $x$ values and assess the validity of the original query. Lastly, our evaluation uses symbolic mathematics to manipulate algebraic expressions. This involves defining symbols and performing operations like factorization, simplification, and algebraic manipulation. The symbols are placeholders for any value, enabling the definition of general expressions without specifying their values upfront.
@@
We evaluate the capabilities of models to orchestrate a multi-step generative process and evaluate a set of tasks. Models need to direct sub-processes and associate computational results from and to Symbol nodes, and maintain relationships between these nodes. Given that the field is currently at an early stage in developing even sequential schedulers for LLM-based planning systems, our initial evaluations will be confined to sequential execution only. In Algorithm 1, we detail our evaluation protocol, which is designed not only to analyze and score a series of instructions but also to provide a structured basis for recording these processes. We also note that our evaluation protocol is generally formulated, which allows the application of non-sequential plan schedulers. In Figure 4 we illustrate a step-wise evaluation of the contextual computational graph, in which the NeSy engine is processing conditioned on the current execution context and producing the next symbol prediction.
@@
1:NeSy engine: $\mathcal{V}:\mathcal{S}\rightarrow\mathcal{S}$, where $\mathcal{S}=\bigcup_{n\geq 0}\mathbb{L}^{n}$ is the set of all possible strings formed by concatenating characters from a character set $\mathbb{L}$, embedding engine $\mathcal{E}:\mathcal{S}\rightarrow\mathcal{H}\subset\mathbb{R}^{d}$, symbols $\{x_{0},x^{*},y^{*}\}\subset\mathcal{S}$, with $x_{0}$ as the initial instruction, $x^{*}$ as the payload resulted from executing $\mathcal{V}$, $y^{*}$ as the reference, and *** acting as a placeholder for $\mathcal{P},\mathcal{T},\mathcal{C}$, capabilities $\mathcal{C}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\ldots\}$, where each $\mathcal{F}_{i}$ represents a specific functional role within the system, plan $\mathcal{P}\subset\mathcal{S}$, task $\mathcal{T}\in\mathcal{P}$, memory buffer $\mathcal{M}\subset\mathcal{S}$, a scoring function $\tilde{s}:\mathcal{H}\times\mathcal{H}\rightarrow[0,1]$, a scheduler $\mathcal{Q}$, an aggregator $\mathcal{A}$, and score variables $\{s\}\in[0,1]$.
@@
3:$\mathcal{V},\mathcal{E},\mathcal{Q},\mathcal{C},y^{\mathcal{P}}\leftarrow$ Init($\cdot$) $\triangleright$ Initialize the engines, the scheduler, the capabilities, and expected plan.
4:$\mathcal{M}\leftarrow\emptyset,\mathcal{A}\leftarrow\emptyset$triangleright Initialize the memory buffer and aggregator.
5:$x^{\mathcal{P}}\leftarrow$ GeneratePlan($x_{0},\mathcal{V}$) $\triangleright$ Generate a plan starting from the initial instruction.
@@
8:while $\mathcal{P}\neq\emptyset$ do $\triangleright$ Run until all tasks have been dequed.
9:     $\mathcal{T},y^{\mathcal{C}},y^{\mathcal{T}}\leftarrow$ Schedule($\mathcal{M},\mathcal{Q}$) $\triangleright$ Select the next task according to the scheduler and progression.
10:     $\mathcal{F}_{i}\leftarrow$ Identify($\mathcal{T},\mathcal{C},\mathcal{V}$) $\triangleright$ Identify the task-related capability $\mathcal{F}_{i}$ using $\mathcal{V}$.
11:     $x^{\mathcal{C}},x^{\mathcal{T}}\leftarrow$ Execute($\mathcal{T},\mathcal{F}_{i}$) $\triangleright$ Execute $\mathcal{T}$ using capability $\mathcal{F}_{i}$ and assign the results $x^{\mathcal{C}},x^{\mathcal{T}}$.
@@
13:     $\mathcal{P},\mathcal{M}\leftarrow$ Update($\mathcal{T},\mathcal{P},\mathcal{M},\mathcal{A}$) $\triangleright$ Update the plan and task progression.
@@
16:return $s$triangleright Return the aggregated score of the plan execution.
@@
Algorithm 1 This algorithm demonstrates the implementation of our Vector Embedding for Relational Trajectory Evaluation through Cross-similarity (VERTEX) protocol. We start by initializing the NeSy engine $\mathcal{V}$, the embedding engine $\mathcal{E}$, the scheduler $\mathcal{Q}$, and a set of capabilities $\mathcal{C}$. The initial instruction $x_{0}$ is used to generate a plan $x^{\mathcal{P}}$ through $\mathcal{V}$. The plan and its expected outcome $y^{\mathcal{P}}$ are embedded, and their similarity is scored and aggregated. The plan is then unfolded into actionable tasks. Each task $\mathcal{T}$ is selected and executed using the appropriate capability $\mathcal{C}$, resulting in the capability and task results $x^{\mathcal{C}},x^{\mathcal{T}}$, and expected outcomes $y^{\mathcal{C}},y^{\mathcal{T}}$ updated in the memory buffer $\mathcal{M}$. The process continues, with each task’s result being embedded, scored, and aggregated until the plan is complete. The final aggregated score $s$ is returned, reflecting the overall effectiveness of the plan execution.
Benchmarks GPT-4 GPT-3.5 Gemini-Pro LlaMA 2 Mistral Zephyr Random Associations 0.93 0.61 1.00 0.41 0.49 0.28 0.01 Modality 0.83 0.77 0.75 0.47 0.39 0.43 0.07 Code 0.78 0.58 0.53 0.10 0.19 0.13 0.00 Logic 0.41 0.44 0.39 0.44 0.46 0.43 0.00 Graphs 0.36 0.31 0.21 0.05 0.06 0.08 0.00 Total 0.66 0.54 0.58 0.30 0.32 0.27 0.02

### v3

@@ 6 Performance Measure
One of the challenges when creating multi-step generative processes with LLMs as part of NeSy engines relies on model evaluation and handling irrelevant predictions. The naïve assessment that measures only task succession would score all models to zero and render them as unusable. Even if models follow instructions and produce parts of the expected solution, we regularly observe that they — especially open-source models — append a continuation of task irrelevant predictions. Such predictions result in failure modes when applying conditions and validations, and halt any multi-step procedure. Our solution is an evaluation protocol that refines the performance measurement, allowing for more nuanced diagnostics and the possibility of continuing the evaluation despite intermediate failures. To derive our quality measure, we borrow ideas from the utilization of the Fréchet distance for generative processes (Heusel et al., 2017).
@@
We generate trajectories through a NeSy sequential process that creates a trajectory of distributions $\mathbb{P}$ over multiple iterations of generative nodes. Each node in the process can be aligned to a reference distribution, which marks the desired behavior. To quantify the validity of the generated trajectories, we measure the total distance between the generated and reference data distribution along the path trajectory. We therefore adopt a cumulative measure capable of taking into account the entire generative trajectory. In theory, this process would entail calculating the path integral over the latent space representations for models, cumulating the Fréchet distances (Dowson & Landau, 1982) traversed along these trajectories:
@@
However, this approach is computationally intractable for large-scale problems, and requires access to latent representations, which — especially in the context of LLMs — is not always given. For computational feasibility, we introduce an approximation that measures the embedding distances over the path trajectories through an auxiliary embedding model, based on prior work on distribution regression (Szabó et al., 2016). The embedding model maps the symbolic representations into a RKHS, such that we can apply a kernel mean embedding function to measure their respective distances (You et al., 2019; Dinu et al., 2023). We assess the distance through the mean embeddings w.r.t. to a kernel function $K(\cdot,\cdot)$ of the samples $\mathbf{e}^{t}_{x}\sim\nu^{t}_{\text{gen}}\in\mathbb{P}_{\text{gen}}$ and $\mathbf{e}^{t}_{y}\sim\nu^{t}_{\text{ref}}\in\mathbb{P}_{\text{ref}}$ produced by the generated data distribution and a reference data distribution respectively. We denote by $\mu_{\mathbf{e}^{t}_{x}}$, $\mu_{\mathbf{e}^{t}_{y}}$ the mean embeddings associated to the respective samples, i.e. $\mu_{\mathbf{e}^{t}_{x}}(z)=\frac{1}{n}\sum_{i=1}^{n}K(x_{i}^{t},z)$ in case $\mathbf{e}^{t}_{x}=(x_{i}^{t})_{i=1}^{n}$ is a sample of size $n$ of the respective mean embeddings. To compute the similarity between the embeddings of the generated and reference distributions, we evaluate the associated maximum mean discrepancy $\text{MMD}^{2}(\mu_{\mathbf{e}^{t}_{x}},\mu_{\mathbf{e}^{t}_{y}})$ (Gretton et al., 2012) and then, as before for the Fréchet distances, we integrate over $t$:
@@
In empirical evaluations, however, we care about normalized values for ease of interpretation. We therefore analyze the properties of the MMD and derive a similarity score, which follows the same statistical principles as the MMD, and is bound between $[0,1]$. We concluded that we can utilize only the MMD cross terms to evaluate the similarities. See Appendix Section B for more details. For our comparisons as referenced in Figure 6 we therefore denote the similarities rather than distances. We then come to the following formulation and refer to our empirical measure as the ”Vector Embedding for Relational Trajectory Evaluation through Cross-similarity”, or VERTEX score for short:
@@
We approximate the integral across time steps through Monte Carlo approximation. The introduced normalization constants denote the similarities to a random sequence $z_{\text{rand}}$, which functions as a baseline subtraction to recenter our results, and a given reference score to rescale w.r.t. to scores obtained from comparing related solutions $z$. Min-max scaling ensures the final measure is bounded between $[0,1]$. This process reflects properties such as Hölder continuity that bounds the kernel function within certain limits. To compute the embeddings, we utilize the embedding model all-mpnet-base-v2 (Song et al., 2020), due to its widespread availability, and its balance between speed and quality. As a similarity measure, we select a Gaussian kernel following our derivation from the Appendix Section B. In our implementations, we also explore other kernels, including preliminary experiments with cosine similarity. We also note that one can integrate Bernoulli distributed trials into our score, with 00 values representing failure modes and values of $1$ being successes. Furthermore, if we relax our definition, we can integrate other similarity measures which are bound between $[0,1]$, which then reflect on domain-specific attributions, i.e. including a similarity measure tailored towards capturing the nuances between two sub-structures of abstract syntax tree.
7 Evaluation
We introduce a benchmark that evaluates multi-step generative processes as complex workflows. Our benchmark consists of five different evaluation categories, and uses the VERTEX score to measure the capabilities of an LLM to solve tasks from each category. The five categories of our benchmark are divided into three basic evaluations and two advanced categories that combine different basic capabilities. The three basic categories are (i) associative prediction which measures a models proficiency in understanding associations between symbols, (ii) multi-modal binding where we evaluate the capability to employ tools and operate on different modalities, and (iii) program synthesis for measuring a models proficiency in generating consistent code and executing subroutines. The two advanced benchmark categories are (iv) logic, for evaluating logic-based components and (v) computational graphs where complex workflows need to be processed, evaluating all aforementioned capabilities. For our evaluation we focus on the GPT family (Brown et al., 2020) of models, specifically GPT-3.5 Turbo (revision 1106) and GPT-4 Turbo (revision 1106) as they are the most proficient models to date; Gemini-Pro (Google, 2023) as the best performing model available through API from Google; LLaMA2-Chat 13B (Touvron et al., 2023), LLaMA3-Chat 8B and LLaMA3-Chat 70B from Meta represent open-source LLMs. Finally, Mistral 7B (Jiang et al., 2023) and Zephyr 7B (Tunstall et al., 2023) serve as baselines for revised and fine-tuned open-source models. The open-source models Mistral, Zephyr, and smaller LLaMA variants are estimated to have roughly equivalent parameter counts compared to GPT-3.5 Turbo and Gemini-Pro. All our experiments require a context size smaller or equal to $4096$ to enable the comparisons among the in-context capabilities across model architectures. For the LLaMA models, we utilize the chat versions since they are specifically tuned to follow instructions.
Associative Prediction
We evaluate a model’s proficiency to follow simple and complex instructions and associations with zero- and few-shot examples. Therefore, we evaluate the proficiency in applying our operators between Symbol types. We defined a total of 15 tasks involving in-context associations between two Symbol instances. SymbolicAI’s overloaded operators rely on predefined pseudo-grammars, as described in Section 4, that augment the operators with few-shot examples. For instance, the overloaded operator + utilized between two Symbol instances provides few-shot examples how to resolve additions with various data types. Consequently, we can now test if the models can solve the addition between Symbol("two hundred and thirty four") and Symbol(7000). See Appendix Section F.1 for more details.
Multi-modal Binding
We perform transformations between multiple modalities through language-based representations. Therefore, we need to evaluate the model’s proficiency in tool utilization, classification and routing of requests to relevant modules. We define a multi-modal Expression to detect the category of a task based on its content and to forward the task to the appropriate tool. The expression creates interfaces to tools like WolframAlpha for mathematical expressions, Selenium for website content scraping, SerpApi for search queries, and APILayer for optical character recognition. Each of the five tests aims to evaluate the appropriate handling of a specific type of input by the multi-modal Expression type, such as processing a website URL for scraping, interpreting a search engine query, testing if two vectors are linearly independent, comparing large numbers, and extracting text from an image. See Appendix Section F.2 for more details.
Program Synthesis
We evaluate executable code with and without concepts from retrieval augmented generation, model-driven development, and experiment with self-generated instructions by creating self-referential expressions. We designed three separate tests related to program synthesis, where each task assesses the ability of the models to generate and execute code based on natural language instructions or provided templates:
1) The first task involves reading a LaTeX table template and data, then generating a function to populate the table with the given data.
2) The second task tests the automatic code generation for API calls by fetching data from a specified URL and extracting specific information from the retrieved content.
3) The third task evaluates the ability to construct a custom Expression that processes a Symbol through a specific Function component from the SymbolicAI package.
Each of the three tests follows a similar pattern, where the generated code is scored based on its similarity to valid references and normalized with random samples. See Appendix Section F.3 for more details.
@@ 7 Evaluation
To evaluation the capabilities for logical reasoning of models, we condition them to create a sequence of expressions as self-contained components, and refer to higher-order logic for their assessment. Based on the underlying type theory originating from Whitehead & Russell (1925–1927), we evaluate a models’ capability to resolve statements in the form of there exists x such that x fulfills y. Such quantifiers define the standard semantics of expressions, where their meaning is given by a semantic function. A semantic function maps a term from an abstract definition to a point in a domain, which is an interpretation of the term’s type and value. Therefore, these functions operate on types and values of expressions, and relations thereof. Subsequently, NeSy engines can formulate and evaluate at inference time logic-based instructions through Lisp, Prolog, or Mathematica (McCarthy, 1959; Colmerauer & Roussel, 1993; Chen et al., 1993; Inc., 2022), or leverage solvers such as Z3 (Moura & Bjørner, 2008). Therefore, the result of a natural language statement when evaluated by a NeSy engine can be interpreted by any expert system which defines the corresponding semantic functions and process them either in a symbolic (Feigenbaum et al., 1965; Gamble et al., 1994), differentiable (Veličković & Blundell, 2021; Ibarz et al., 2022), or hybrid manner (Kuncicky et al., 1991).
We evaluate how proficient models are at interpreting custom DSLs and define expression statements. DSLs are designed to express logical relations and operations in a structured format, and supports human-readable and machine-interpretable formulations. The following example illustrates such relationships by translating a natural language statement into an expression statement, as follows:
@@
A DSL may enforce the usage of $\text{HAS}(\cdot)$, $\text{IS}(\cdot)$, etc. and may condition an LLM to produce the following expressions:
@@
These are then utilized to define the following logical expression:
@@
An automated theorem prover can now evaluate this statement for all values of $x$ and assess the validity of the original query. Lastly, our evaluation uses symbolic mathematics to manipulate algebraic expressions. This involves defining symbols and performing operations like factorization, simplification, and algebraic manipulation. The symbols are placeholders for any value, enabling the definition of general expressions without specifying their values upfront.
We designed six tests to assess the logical capabilities of the candidate models and group them as follows. See Appendix Section F.4 for more details.
1) We utilize the Python library SymPy for symbolic mathematics to create the mathematical expression $ax+bx-cx-ay-by+cy+d$. The task for the model is then to factorize the expression and extract all unique symbols as a list.
2) Three tasks evaluate a models’ capability to resolve the logical operations AND, OR, and XOR. For instance, the test for logical AND combines the symbols Symbol("The horn only sounds on Sundays") and Symbol("I hear the horn") and compares the answer against the human-generated references ”The horn only sounds on Sundays and I hear the horn.” and ”Since I hear the horn it is Sunday.” Since there is a large number of possible solutions, there is high variability in the solution space. Each model might prefer a different solution.
3) For another task we use a custom Expression that defines a DSL syntax and semantic structure. We use this Expression to extract higher-order logic expressions from a natural language statement, namely the puzzle ’Who is Jay’s brother?’444 Bob has two sons, John and Jay. Jay has one brother and father. The father has two sons. Jay’s brother has a brother and a father. Who is Jay’s brother?, that preserves the original relationships.
4) For the final task, we again use the puzzle ’Who is Jay’s brother?’ to evaluate a models’ capability for complex conversions. We use the Z3 theorem prover (Moura & Bjørner, 2008) to solve the ’Who is Jay’s brother’ puzzle conditioned on the Z3 solvers’ solution to Einsteins’ famous puzzle ’Who owns the fish?’. The task involves an indirect translation from natural language to executable code by the Z3 solver; the solution to Einstein’s puzzle acts as a form of self-contained ”documentation” for how the Z3 solver should be utilized. The test constructs a template, which includes the task instructions, puzzle statement, and reference to the Einstein’s puzzle solution. The models are then asked to analyze the given problem and solution format and create a Python function with Z3 syntax that can solve the ’Who is Jay’s brother?’ puzzle. The dynamically generated code is executed within the test environment utilizing Python’s exec function. We check the access to the Z3 solver and run the generated solve_puzzle function supposed to contain the logic to solve the puzzle. Once executed, the assembled Z3 logical clauses are processed by the solver, which verifies that the set of constraints is satisfiable. If so, the model generated by the solver is queried for the puzzle’s solution and scored using our VERTEX score.
@@
We evaluate the capabilities of models to orchestrate a multi-step generative process and evaluate a set of tasks. Models need to direct sub-processes and associate computational results from and to Symbol nodes, and maintain relationships between these nodes, which we refer to as a computational graph as shown in Figure 5. In a computational graph, the VERTEX score compares the results produced by a generative model at each node against samples obtained from a reference distribution, usually modeled by sampling from multiple valid references. We also account for randomness through predefined random samples for normalizing the result. Our reference to hierarchical computational graphs stems from the fact that we operate on multiple levels. On a higher level of abstraction we are able to perform planning, sub-task scheduling, and define operational instructions. On a lower level of abstraction, we execute these plans based on the defined instructions and data, which can also span generative processes that produce new information.
Figure 5: We illustrate the hierarchical computational graph for the Paper expression. Each node represents an instance of an expression with distinct properties and behaviors, such as file sourcing, generative process, tool utilization, or transformation operation. The edges denote the reference relationships between expressions and indicate the flow of information. The blue highlighted nodes mark the main sequence nodes of expressions utilized to create parts of the paper draft, such as Method section, RelatedWork section, Abstract section, and so on. Each generative node is used for evaluating the VERTEX score. None-generative nodes such as search engine results are not evaluated, and we assume to obtain ground-truth values.
Given that the field is currently at an early stage in developing even sequential schedulers for LLM-based planning systems, our evaluations will be confined to sequential execution only. We introduce two tests designed to evaluate multi-step generative processes:
1) We simulate and evaluate the process of writing a research paper draft based on a predefined hierarchical computational graph that focuses on the content output of the computational graph rather than planning and scheduling functionality. See Appendix Section F.5 for more details.
2) We test the VERTEX Protocol as defined in Algorithm 1, which represents our general method for evaluating multi-step generative processes. We create a self-contained test scenario to illustrate an end-to-end evaluation and as a go-to reference for how our protocol can be deployed in a realistic environment. Our evaluation protocol is not only designed to analyze and score a series of instructions, but also to provide a structured basis for recording these processes. Furthermore, we note that our evaluation protocol is generally formulated, which allows the application of non-sequential planning and scheduling.
@@
1:NeSy engine $\mathcal{V}$ as an LLM, embedding engine $\mathcal{E}:\Sigma\rightarrow\mathcal{H}\subset\mathbb{R}^{d}$, symbols $\{x_{0},x^{*},y^{*}\}\subset\Sigma$, with $x_{0}$ as the initial instruction, $x^{*}$ as the payload resulted from executing $\mathcal{V}$, $y^{*}$ as the reference, and * acting as a placeholder for $\mathcal{P},\mathcal{T},\mathcal{C}$, capabilities $\mathcal{C}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\ldots\}$, where each $\mathcal{F}_{i}$ represents a specific functional role within the system, plan $\mathcal{P}\subset\Sigma$, task $\mathcal{T}\in\mathcal{P}$, memory buffer $\mathcal{M}\subset\Sigma$, a scoring function $\tilde{s}:\mathcal{H}\times\mathcal{H}\rightarrow[0,1]$, a scheduler $\mathcal{Q}$, an aggregator $\mathcal{A}$, and score variables $\{s\}\in[0,1]$.
@@
3:$\mathcal{V},\mathcal{E},\mathcal{Q},\mathcal{C},y^{\mathcal{P}}\leftarrow$ Init($\cdot$) $\triangleright$ Initialize engines, scheduler, capabilities, expected plan.
4:$\mathcal{M}\leftarrow\emptyset,\mathcal{A}\leftarrow\emptyset$triangleright Initialize memory buffer and aggregator.
5:$x^{\mathcal{P}}\leftarrow$ GeneratePlan($x_{0},\mathcal{V}$) $\triangleright$ $\mathcal{V}$ generates plan based on initial instruction.
@@
8:while $\mathcal{P}\neq\emptyset$ do $\triangleright$ Run until list of tasks is exhausted.
9:     $\mathcal{T},y^{\mathcal{C}},y^{\mathcal{T}}\leftarrow$ Select($\mathcal{M},\mathcal{V}$) $\triangleright$ $\mathcal{V}$ selects next task based on task progression.
10:     $\mathcal{F}_{i}\leftarrow$ Identify($\mathcal{T},\mathcal{C},\mathcal{V}$) $\triangleright$ $\mathcal{V}$ identifies task-related capability $\mathcal{F}_{i}$.
11:     $x^{\mathcal{C}},x^{\mathcal{T}}\leftarrow$ Execute($\mathcal{T},\mathcal{F}_{i},\mathcal{Q}$) $\triangleright$ $\mathcal{Q}$ executes $\mathcal{T}$ with capability $\mathcal{F}_{i}$ and assign results $x^{\mathcal{C}},x^{\mathcal{T}}$.
@@
13:     $\mathcal{P},\mathcal{M}\leftarrow$ Update($\mathcal{T},\mathcal{P},\mathcal{M},\mathcal{Q}$) $\triangleright$ $\mathcal{Q}$ updates plan and progression.
@@
16:return $s$triangleright Return aggregated score of plan execution.
@@
Algorithm 1 This algorithm defines the pseudocode of our VERTEX protocol with our respective VERTEX score as a scoring criteria. We start by initializing the NeSy engine $\mathcal{V}$, the embedding engine $\mathcal{E}$, the scheduler $\mathcal{Q}$, and a set of capabilities $\mathcal{C}$. The initial instruction $x_{0}$ is utilized to generate a plan $x^{\mathcal{P}}$ through $\mathcal{V}$. The plan and its expected outcome $y^{\mathcal{P}}$ are embedded, and their similarity is scored according to our VERTEX score and aggregated. The plan is then unfolded into actionable tasks. Each task $\mathcal{T}$ is selected and executed with the appropriate capability $\mathcal{C}$, resulting in the capability and task results $x^{\mathcal{C}},x^{\mathcal{T}}$, and expected outcomes $y^{\mathcal{C}},y^{\mathcal{T}}$ updated in the memory buffer $\mathcal{M}$. The process continues, with each task’s result being embedded, scored, and aggregated until the plan is complete. The final aggregated score $s$ is returned, reflecting the overall effectiveness of the plan execution.
We start with a high-level workflow description which consists of a list of tasks and optionally their respective sub-tasks; we refer to this as the plan $\mathcal{P}$. To perform the experiment, we utilize an expected plan $y^{\mathcal{P}}$ which was handcrafted for this evaluation. The expected plan is a queue of predefined tasks (in a particular order) that the system should follow to achieve the goal. The goal statement defines the end objective that the workflow aims to accomplish. We also have a set of plans similar to the expected plan, which are trajectories in the solution space, as well as the plan $x^{\mathcal{P}}$ that the LLM generates utilizing the GeneratePlan call for a specific seed. We score the predicted plan against the expected plan and the trajectories, then we continue to the next phase in which we utilize the expected plan to execute the tasks. At each step, the LLM will receive in its context the goal, the tasks, the current progress, and a query asking for the next task to execute; we refer to this as the memory buffer $\mathcal{M}$. If the LLM is not able to predict the next task, it will return a failure, and the expected plan will be utilized to execute the next task. The LLM has access to a predefined set of capabilities $\mathcal{C}$, specifically WolframAlpha, SerpApi, Selenium, and the LLM itself, which also represents our self-referential structure. We keep executing tasks until the queue is exhausted, and at each step, we utilize the Evaluate call to measure the performance of the LLM with our VERTEX score. The scheduler class $\mathcal{Q}$ oversees the execution of the test workflow. It takes the setup configuration and orchestrates the linear execution of tasks, utilizing the expected plan as a reference. It maintains a pool of tasks to be executed and updates progress as tasks are completed. The UnfoldPlan call is a method of the scheduler class $\mathcal{Q}$. The method calls itself recursively until there are no tasks left. The Select call is responsible for determining which task to execute next from a pool of remaining tasks. It utilizes the LLM through self-reflection (Shinn et al., 2023) to choose the most suitable next task based on a template that gets progressively updated in the memory buffer $\mathcal{M}$ by the Update call. The Identify call uses self-reflection and similarity scoring to determine the best interface based on the task at hand, then passes the interface to the Execute call to execute the task. Lastly, the test ends with the Finalize call, which provides an aggregated assessment of the model’s ability to manage and execute the workflow.
In Figure 6 we conclude with our evaluation and compute the cumulative score for all described evaluation categories and in the next section we discuss how to interpret the results of our framework.